Three nano coulombs (nC) are placed at a seperate corner of a square with each side measuring 3cm. The two charges located at opposite vertices are negative, and other given charge is positive. Calculate the force experienced by a fourth 3nC when place at the missing corner. What is it orbital period and what its radial acceleration?

1
Expert's answer
2022-04-21T13:01:34-0400

"F=\frac{kq^2}{(1+(\sqrt 2)^2+1)r^2}=32~kN."